sara invests $3000 in a mutual fund. on average the mutual fund earns 9%/yr compounded annually. how much interest can sara expect to earn after 10 yr
Answer by Theo(13342)

f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

in your problem.

p = 3000
r = .09 per year.
n = 10 years.

solve for f to get f = 3000 * (1 + .09) ^ 10 = 7102.091024

i = f - p

i equals interest.
f = future value
p = present value

in your problem:

i = 7102.091024 - 3000 = 4102.091024

you can combine the two formulas into one as shown below:

i = p * (1 + r) ^ n - p